OpenStudy (anonymous):

Use implicit differentiation to find (dy/dx) of 4xy+5ln(y)=28 at the point (7,1).

OpenStudy (abb0t):

\(4x \frac{dy}{dx}(y)+y \frac{dy}{dx}(x)+5[\frac{dy}{dx}(ln(y))]=0\) \(\Rightarrow\) \(4xy'+y+5\frac{1}{y}y'=0\) NOW, solve for y' Note: derivative of a constant is zero, which is why it's equal to zero. and also, NOTE: \(y' = \frac{dy}{dx}\) I was just using short notation

OpenStudy (dan815):

abb0t! dont give up on my question plz!! ur the only one who has attempted so far

OpenStudy (dan815):